Cameron Corp. purchased a mine on January 1, 2018, for $529,000, which is estimated to contain 40,000

tons of iron ore. There is no residual value. The business extracted and sold 14,500

tons of ore in 2018 and 4,800 tons of ore in 2019. What is the depletion expense for 2019? (Round any intermediate calculations to two decimal places, and your final answer to the nearest dollar.)

A) 273,861

B)255339

C)63504

D)191835
Answer #1

Depletion per unit=(Cost-Residual value)/Useful Life

=(529,000/40,000)=$13.23 per ton

Hence depletion expense for 2019=(13.23*4800)

=$63504